OpenStudy (anonymous):

Question: As a salesperson at Snow Sports, Micheal receives a monthly base pay plus commission on all that he sells. If he sells $600 worth of merchandise in one month , he is paid $344. If he sells $1000 worth of merchandise, he is paid $440. Find Michael's salary when he sells $2,200 worth of merchandise.

OpenStudy (anonymous):

thats an interesting question actually

OpenStudy (anonymous):

let x= commission (%) y=salary then: 600x+y=344 1000x+y=440 solve system & find the commission (x as %) & salary (y)

OpenStudy (anonymous):

also u have multiple accounts, this question was asked last night :|

OpenStudy (anonymous):

subtract first equation from second you get: 400x=96 x=0.24 (24%) then y=200 (salary)

myininaya (myininaya):

oops 1*1000s+2*600s*

OpenStudy (anonymous):

ok maybe here. you want the equation for the line between (600,344) and (1000,440) the slope of this line is \[\frac{440-344}{1000-600}=\frac{96}{400}=.24\] an and then use \[y-344=.24(x-600)\] \[y=.24x+200\] for the equation of the line. plug in 2200 for x to get your answer

myininaya (myininaya):

i think this can be made simpler

OpenStudy (anonymous):

then salary when he sales $2200: 2200*.24 +200=728 answer: $728

OpenStudy (anonymous):

btw the line says you get .24 for every dollar in sales and 200 base salary. i get $728

myininaya (myininaya):

i guess it cannot be made simpler

OpenStudy (anonymous):

@ myininaya yes simpler but you still have to find the amount per sale which is basically the slope yes?

myininaya (myininaya):

yes i guess

myininaya (myininaya):

lol

OpenStudy (anonymous):

and you still have to find the base salary aka y- intercept so there is no way around getting the equation for the line

OpenStudy (anonymous):

you can find the y - intercept by doing something else, like plugging in a known value instead of using point - slope, but it make no difference in that you still need to get \[y=mx+b\] in some way

OpenStudy (anonymous):

Okay I think I get it now. Thanks guys :)